OpenStudy (anonymous):

Check my answer please? What is the exact value for the expression the sqrt(171) − sqrt(19) + sqrt(76)? Simplify if possible. My answer 4 sqrt(19)

OpenStudy (anonymous):

4 * sqrt(19) is correct

OpenStudy (anonymous):

yes Correct!

OpenStudy (e.mccormick):

Yah, it looks like you did something like this: \(\sqrt{171}-\sqrt{19}+\sqrt{76}\) \(\sqrt{9\cdot 19}-\sqrt{19}+\sqrt{4\cdot 19}\) \(3\sqrt{19}-\sqrt{19}+2\sqrt{19}\) \(4\sqrt{19}\) Which is perfect.

OpenStudy (anonymous):

Thank you guys!!!